OpenStudy (anonymous):

Which of the following is the reciprocal of x^2 – 2x – 8?

OpenStudy (anonymous):

is it f(x)=x^2 - 2x - 8?

OpenStudy (anonymous):

this is not one of the answer choices

OpenStudy (anonymous):

i know, but is the original question that?

OpenStudy (anonymous):

what are your answer choices?

OpenStudy (anonymous):

the original question was copied and pasted so this is the original question yes

OpenStudy (anonymous):

A. (x + 2)(x – 4) B. 1/(x + 2)(x – 4) C. 1/(x – 8) D. 1/(x + 2)

OpenStudy (anonymous):

what are the answer choices? and you start off by switching the y for the x values so instead of: y=x^2 - 2x - 8, it will be: x=y^2 - 2y - 8, then you solve for y

OpenStudy (anonymous):

I just posted my answer choices

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

i'm still working it out hang on

OpenStudy (anonymous):

I got B.

OpenStudy (anonymous):

yep, good job